Sol Summary – January 02nd

  

 Sol: 11

Summary Title: Mars Stars

Author's name: Pat Pesa

Mission Status: Can't wait until we recieve our Hollywood Walk of Fame Star!

Sol Activity Summary: Our friendly neighboorhood Martian film crew came to document our daily life and some beautiful Mars landscapes. They followed us on our two EVAs (being very respectable of allowing us to still acomplish our work), and after we shared cookies, mission patches, and stories of travel.

Look Ahead Plan: Tomorrow we will perform our last EVA, and then start cleaning and preparing for the end of our Mission.

Anomalies in work: None

Weather: Partly Cloudy , around 34 degrees

Crew physical status: Healthy

EVA: EVA #10 this morning went to marble ritual for seismic tests. EVA #11 this afternoon went to the same location as EVA 8 of yesterday, Galileo road.

Reports to be filed: sol summary, commander report, operations report, greenhab report, journalist report, EVA report, EVA request, science report

Support Requested: none

Pat Pesa

Geologist, MDRS 218

Journalist Report – January 1st

Journalist Report
Ben Durkee, Crew 218 Journalist

Sol 10

The Roaring Twenties: Episode Two

Happy New Year, all! It will be very interesting to see what Earth is like in the new decade when we head back a few days from now. Are there flying cars? Jetpacks? Can we speak Dolphin yet? The
possibilities are endless!

Meanwhile, we are having a blast one planet farther from the sun! We hosted the two MDRS assistant directors and together indulged in a wonderful meal courtesy of Chef Cesare. Streamers were popped, games were played, raucous laughter was shared, but most importantly we partook in some juicy controversy.

Favorite movies were discussed (good and bad), alongside whether or not water is wet, which way the toilet paper roll should hang, how the “g” in “gif” is pronounced, and many more that got progressively stranger as midnight approached. As far as I’m concerned, nothing brings people together quite like witty banter and inconsequential conflict. We entered the new year with stronger friendships, streamer debris all over the Hab, mild headaches from the paper trumpets, and smiles on our faces.

The only downside of New Year’s Night is waking up at the regular time the next morning. Our wake-up train chugged a little slower today as Sol 10 came to fruition. Double digits! Two calendar-based milestones today; the stars truly aligned. We made up for our lack of energy with a breakfast of biscuits and gravy, the recipe for which was sent from Earth by a friendly Mars veteran.

After devouring the edible southern hospitality, our hive was back to full operational capacity and buzzing to prepare Pat and Cesare for our 8th EVA of the mission. They finagled their EVA suits into our loyal rovers and took them beyond range of the Hab’s communication capability. The sun crawled across the cloudy Martian sky and soon enough they were back with plentiful geological data and samples. We were gracious enough to let them and their new pet rocks back in the airlock, which proved to be a mistake as the coveted Nutella supply seemed to drastically deplete with their re-entry.

To recharge the rovers and ourselves, we all ate lunch and recapped the EVA. After a short break, it was time for the second EVA of the day, featuring LuzMa, Shefali, and yours truly. We strictly adhered to Shefali’s new EVA checklist (likely inspired by the infamous shoe incident) and were outside in record time. We saddled up our rovers only to find that they were still tired from lugging the other guys and their equipment.

Unfortunately the recent sustained Martian cold has been hugely detrimental to the rover batteries. They now discharge rapidly and inconsistently, and recharge at a snail’s pace. Too bad it couldn’t have been vice versa – I guess that old adage about batteries lasting longer after refrigeration might not be valid on Mars. Myth busted. We rolled with the punches and turned our roving EVA into a walking EVA around the Hab. We gathered what data we could (redundant data is far better than no data!) and enjoyed the snowy walk back home together. Although it was a bit disappointing, Mars is a fickle creature and we must do the best we can in spite of variables beyond our control.

Speaking of which, the consistent negative temperatures are still preventing our pipes from thawing. We now consider our water assembly line a charming team-building exercise, so we put on some tunes and got to bucketing. Faster than ever, we had the loft tank back to full. Plenty of water for drinking, dinner, dishes, and desperately – showers. We may or may not have a filming crew flying in from Earth tomorrow, and we gotta look (and smell) good for the cameras!

As crew journalist, these guys are kinda threatening my job security. But on the other hand, this could be my big break!

Operations Report – January 1st

Crew 218 Operations Report 1-Jan-20

SOL: 10

Name of person filing report: Luz Maria Agudelo Urrego

Non-nominal systems: NA

Generator: run

Hours run: 12

From what time last night: 1930

To what time this morning: 0730

List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before the generator is run at night: 62%

Diesel Reading – 74%

Station Propane Reading – 65%

Ethanol Free Gasoline: N/A

Water (loft tank) (gal): 55

Water Meter (units): 0147212.1

Water (static tank) (gal): 270

Static to Loft Pump used – No

Water in Green Hab (gal): 84

Water in Science Dome (gal): 0

Toilet tank emptied: Yes

Deimos rover used: Still in the workshop

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: Assigned to director

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: EVA 8

Hours: 121.5

Beginning charge: (Before EVA): 100%

Ending charge: (On return from EVA, before recharging): 29%

Currently charging: Yes

Spirit rover used: EVA 9 – Rover was not used

Hours: 121.5

Beginning charge: (Before EVA):

Ending charge: (On return from EVA, before recharging):

Currently charging: Yes

Opportunity rover used: Still in the workshop

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: EVA 8

Hours: 126.9

Beginning charge: 100%

Ending charge: (On return from EVA, before charging): 40%

Currently charging: Yes

Curiosity rover used: EVA 9 – Rover was not used

Hours: 126.9

Beginning charge:

Ending charge: (On return from EVA, before charging):

Currently charging: Yes

Notes on rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use: N/A

Oil Added? No

# Hours the ATVs were used today: 0

Notes on ATVs: N/A

HabCar used, and why, where? No

CrewCar used, and why, where? No

General notes and comments: The tablespoon in the Hab kitchen is broken.

Summary of the internet: Nominal

Summary of suits and radios: On the two-piece space simulator Suit N°6, the right air pipe connector sheared.

Summary of Hab operations: The pipe from the static tank through the loft tank is still frozen. The crew filled out the Loft Tank by carrying water buckets from the Static Tank again.

Summary of Science Dome operations: Nominal

Summary of RAM operations: Nominal

Summary of any observatory issues: The fire alarm went on at 1820. The crew replaced the batteries in it.

Summary of health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: NA

Science Report – January 1st

Crew 218 Science Report 01-JAN-20
Crew Science Officer: Dr. Jonathan R. Buzan

1. Decision Making in support of autonomy for crew EVAs:

No EVA: nothing to report.

2. Mars surface weather:

Measurements were collected during today’s EVA.

3. Subsurface structure on Mars:

Seismic reflection mapping instrument was deployed in today’s EVA.

4. Detecting radio signal strength:

Field measurements were collected during today’s EVA.

5. EVA workload analysis:

Survey data was collected for today’s EVAs.

6. EMU (Extravehicular Mobility Unit) ergonomic assessment:

Survey data was collected for today’s EVAs.

7. Environmental Stresses over MDRS habitat and Crew Members and projection over Martian Terrain:

Interior measurements not planned today: nothing to report.

8. Messier and other space objects for outreach:

Cloudy: nothing to report.

9. Reliability and maintenance:

Radio and rover failure procedures evaluated.

10. Medical readings in preparation for future crew-wide project:

The Zephyr bio-monitor was worn all day.

11. Collection of clay, shale, and hematite samples:

No EVA: nothing to report.

12. Media and outreach:

Nothing to report.

Glassware check out: None

A reminder to all crewmembers: There’s a $300 fine for using any glassware material without MDRS Mission support permission.

GreenHab Report – January 1st

Crew 218 GreenHab Report 01-JAN-20

Crew GreenHab Officer: Dr. Jonathan R. Buzan and HSO Shefali Rana

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 23.5°C; 18%

12:15PM

Floor Unit: 20°C

Electronic: 27°C

humidity 16%

3:20PM

Floor Unit: 18°C

Electronic: 22°C

humidity 19%

6:30PM

Floor Unit: 15°C

Electronic: 19.1°C

humidity 20%

Max: 30.1; 20%

Min: 16.8°C; 16%

Hours of supplemental light: Light system 6:00PM-11:00AM.

Daily water usage for crops: 3.7Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – ~84 Gal.

Time(s) of watering for crops:

6:30PM. 3.7 Gal.

SEASONAL PLANTING HISTORY:

Change to crops: None.

OTHER NOTES:

1st sprouts: N/A

Harvest: N/A

EVA Report – January 1st

Crew 218 EVA Report 01-01-2020

EVA #: 8

Author: Cesare Guariniello

Purpose of EVA: Seismometric measurements; EVA and EMU evaluation project; reliability project

Start time: 10:57

End time: 12:59

Narrative: Pat and Cesare drove East on Galileo / Cactus road on a cold, battery-draining day. Once they found a suitable spot in the lowlands they set up the seismometric experiment. Then, they performed a reliability and repair study, evaluating the execution of operations via radio communication. Pat served as the expert, guiding Cesare via radio. The experiment simulated a rover battery failure to be repaired in 10 steps, which included cleaning the area around the seats from ice, putting the map in the glovebox, and reciting poetry. After this, Pat and Cesare performed seismometric readings before heading back to the hab

Destination: Galileo / Cactus Rd

Coordinates (use UTM NAD27 CONUS): 12S 519850, 4252000

EVA Participants: Cesare Guariniello, CMD; Pat Pesa, GEO

Road(s) and routes per MDRS Map: Cow Dung Road and Galileo / Cactus Rd

Mode of travel: Driving, then work in the field and walking

Vehicles you will be using (If applicable): Curiosity and Spirit

EVA #: 9

Author: Cesare Guariniello

Purpose of EVA: Weather observation; EVA and EMU evaluation project; reliability project

Start time: 14:18

End time: 14:54

Narrative: After heading out, the crew found out that the rovers were not fully charged. The commander called for cancellation of the driving, and allowed the crewmembers to perform weather observation and reliability project in the vicinity of the habitat. The crew reached the MDRS sign and simulated a situation of radio failure, with consequent EVA abortion and implementation of protocols for
communication and safety, which included having the crewmember experiencing a radio failure to walk between the two crewmembers with working apparatus.

Destination: Brahe Hwy (intended), vicinity of the hab (actual)

Coordinates (use UTM NAD27 CONUS): 12S 519850, 4252000

EVA Participants: Ben Durkee, JOU; Luz Ma Agudelo, ENG; Shefali Rana, HSO

Road(s) and routes per MDRS Map: vicinity of the habitat

Mode of travel: Walking

Vehicles you will be using (If applicable): None

Cesare Guariniello, Commander

Sol Summary – January 1st

Sol: 10

Summary Title: MARSching Into the Next Year

Author’s name: Pat Pesa

Mission Status: Best Crew of the Decade!

Sol Activity Summary: We once again embarked on two EVAs to continue catching up on our research from the first week. However we still ran into some troubles in the second EVA because we did not allow for enough time so that the Rovers were fully charged and thus the second EVA was limited to walking range of the Hab.

Look Ahead Plan: Tomorrow along with more EVAs we also are expecting some TV crew to document the station and what we do on a daily basis.

Anomalies in work: None

Weather: Partly Cloudy , around 34 degrees

Crew physical status: Healthy

EVA: EVA #8 this morning went to Galileo Road where we performed some more seismic tests and a decision making matrix test. EVA #9 this afternoon was cut short due to Rovers not being fully charged yet, but there was still some research done close to the Hab.

Reports to be filed: sol summary, commander report, operations report, greenhab report, journalist report, EVA report, EVA request, science report

Support Requested: none

Pat Pesa
Geologist, MDRS 218

Commander Report – January 1st

Crew 218 Commander Report 01-01-2020

Sol 10 – Welcome, 2020

In the middle of our last week at MDRS, and after our New Year’s Eve celebration, we woke up at 8.30 for another productive day. After breakfast, an EVA in very cold and cloudy weather reached for the first time in this mission a location distant from the main road, with a good ride and field work off Galileo Rd. Unfortunately, cold weather also means slow battery recharge. The second EVA planned for the day could not therefore use the rovers, and the two research projects associated with it were performed in the vicinity of the habitat, after an engineering EVA. Work is proceeding well, though the commander felt the need for a good nap before working on reports.

Cesare Guariniello, Commander